OpenStudy (anonymous):

A net force of 12.0N north acts on a object for 2.00*10^-3 . calculate the impluse

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

hello there..impulse is that force which is applied on to an object for a very short period of time it is given by J = F * dt where F is the force and dt is the small time interval now F= 12N dt=2*10^-3 calculate J by substituting these values :) and its unit is Ns

OpenStudy (anonymous):

impulse is the integral of the resultant force (F) with respect to time: \[J = \int\limits_{t1}^{t2}F.dt\]